# An accurate mean-field equation for voter model dynamics on scale-free networks

### v1

@@ III Robustness
The robustness of $\alpha$ is verified with the following numerical study. We consider a BA network with preferential attachment parameter $m=3$ and $N=10^{5}$ nodes, and the CNVM with $M=3$ opinions. We sample 200 initial states $x\in[M]^{N}$ using four different strategies to obtain a diverse set of random initial conditions:
@@
    Uniform shares: Sample shares $c$ uniformly at random from the simplex $\{c\in[0,1]^{M}\mid\lVert c\rVert_{1}=1\}$, and assign opinions randomly so that $c$ is matched.
@@
    By degree: Again, sample shares $c$ uniformly. Choose a random order of opinions, assign the first opinion to the $Nc_{1}$ nodes with largest degree, the second opinion to the next $Nc_{2}$ remaining nodes with largest degree, and so on.
@@
(In fact, every initial state would result in approximately the same $\alpha$, as this experiment will show.) For each of the 200 initial states and parameter sets, we computed the trajectories of the weighted shares $\bar{c}(t)$ via simulation and compared them to the mMFE solution $\bar{c}_{\text{{\tiny mMFE}}}(t)$ under identical initial conditions and parameters.
The results of this experiment are shown in Fig. 4. We observe errors below 0.005 for all trajectories, with a median error smaller than 0.003. This confirms that our choice of $\alpha=0.81$ yields a very small error for any choice of initial condition and parameter set in regimes A and B, thereby demonstrating the robustness of the mMFE.
@@
As mentioned before, the proper choice of $\alpha$ depends only on the preferential attachment parameter $m$. Since $m$ determines the minimum node degree, a larger $m$ generally reduces neighbor correlations. Consequently, as $m$ increases, the optimal $\alpha$ also increases. For very large $m$, neighbors are almost independent, implying $\alpha\approx 1$. This behavior is illustrated in Fig. 5, which shows the optimal $\alpha$ computed from (22) for different values of $m$.
It should be noted that the above results—especially regarding the validity of the mMFE and the robustness of $\alpha$—are accurate only for sufficiently large networks. In the examples discussed here, we found that $N\geq 10^{4}$ is a reasonable threshold. For smaller $N$, the mMFE, like other mean-field approximations, loses accuracy.
@@ IV Comparison to classical approximations
Figure 6: The shares $c$ and weighted shares $\bar{c}$ estimated from simulations, in comparison to the mMFE, MFE, and PA. Model parameters from regime A. The left plot shows the trajectories, while the right plot shows the $L_{1}$-error (19) between the mMFE and $\bar{c}$, as well as between MFE and $c$, and PA and $c$.
In this section we compare the mMFE, which is a (M1)(M-1)-dimensional approximation, to other standard low-dimensional moment-closure approximations: the MFE, which also has dimension (M1)(M-1), and the pair-approximation (PA), which has a dimension of (M1)+M(M1)(M-1)+M(M-1). The PA results from closing the moment equations at second order and therefore includes the number of edges between different opinions as additional variables compared to the MFE. For a derivation of the PA, see for example [23]. As Fig. 6 illustrates, the classical approximations (MFE and PA) fail to capture the system behavior on scale-free networks, whereas the mMFE incurs only a very small error. This is not surprising, since the classical approximations assume a homogeneous network structure. The well-known heterogeneous variants of the MFE and PA, or the approximate master equation, are better suited for scale-free networks [24, 4, 15, 25]. However, as these resolve the opinion shares per degree, they involve substantially more variables than the mMFE. For instance, on a BA network with $N=10^{4}$ nodes and $m=4$, there are on average about 100 different node degrees, implying that the heterogeneous MFE has 100 times more variables than our mMFE. Hence, to compare reduced order models with the same degree of efficiency, we only report comparison between the mMFE and the homogeneous MFE and PA.
@@ V Conclusion
In this paper, we derived the mMFE (3), a modified version of the classical mean-field equation (MFE) that describes the dynamics in terms of weighted shares and includes an average “correlation factor” $\alpha$. We then verified that this equation provides an accurate and efficient approximation of the continuous-time noisy voter model on scale-free networks generated by the Barabási–Albert model. It outperforms the standard MFE and pair-approximation (PA), while remaining low-dimensional. We found that the approximation error depends on the underlying parameter regime; in particular, in regimes C and D—corresponding to (partially) noise-free dynamics and/or very dissimilar imitation rates—a time-constant correlation factor $\alpha$ does not exist, causing larger errors.
The inspiration for the mMFE (3) arose from findings in [19], where a computational approach was proposed to identify forecastable variables of dynamical processes on networks. For such processes, reduced modeling is typically guided by theoretical insights—our approach therefore constitutes an exception. Along these lines, it would be highly desirable to gain a theoretical understanding of the observed superior behavior of the mMFE. Furthermore, this work also highlights the importance of data-driven approaches, since many of the insights presented here were supported by computational analysis. These methods could be utilized in future work to derive low-dimensional approximations for other systems that are not yet well understood analytically.
This work in particular could also be extended by refining the mMFE in parameter regimes where its accuracy is limited, for instance by allowing a time-dependent $\alpha$. Moreover, a modified version of the PA that operates on the degree-weighted shares $\bar{c}$ could be explored.
@@ Appendix A Derivation of mMFE for more than two opinions
In the case of $M>2$ opinions, the degree-weighted shares $\bar{c}\in[0,1]^{M}$ are defined via $\bar{c}=(\bar{c}_{a})_{a\in[M]}$,
@@
$\frac{\text{d}}{\text{d}t}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}]$ $=\sum_{b\neq a}\mathbb{P}(x_{i}=b)\ \mathbb{E}[q_{b,a}^{i}\mid x_{i}=b]$ (27)
$\ -\sum_{b\neq a}\mathbb{P}(x_{i}=a)\ \mathbb{E}[q_{a,b}^{i}\mid x_{i}=a].$ (28)

### v3

@@ III Robustness
The robustness of $\alpha$ is verified with the following numerical study. We consider a BA network with preferential attachment parameter $m=3$ and $N=10^{5}$ nodes, and the CNVM with $M=3$ opinions. We sample 200 initial states $\bm{x}\in[M]^{N}$ using four different strategies to obtain a diverse set of random initial conditions:
@@
    Uniform shares: Sample shares $\bm{c}$ uniformly at random from the simplex $\{\bm{c}\in[0,1]^{M}\mid\lVert\bm{c}\rVert_{1}=1\}$, and assign opinions randomly so that $\bm{c}$ is matched.
@@
    By degree: Again, sample shares $\bm{c}$ uniformly. Choose a random order of opinions, assign the first opinion to the $Nc_{1}$ nodes with largest degree, the second opinion to the next $Nc_{2}$ remaining nodes with largest degree, and so on.
@@
(In fact, every initial state would result in approximately the same $\alpha$, as this experiment will show.) For each of the 200 initial states and parameter sets, we computed the trajectories of the weighted shares $\bar{\bm{c}}(t)$ via simulation and compared them to the mMFE solution $\bar{\bm{c}}_{\text{{\tiny mMFE}}}(t)$ under identical initial conditions and parameters.
The results of this experiment are shown in Fig. 4. We observe errors below 0.005 for all trajectories, with a median error smaller than 0.003. This confirms that setting $\alpha=0.81$ yields a very small error for any choice of initial condition and parameter set in regimes A and B, thereby demonstrating the robustness of the mMFE.
@@
As mentioned before, the proper choice of $\alpha$ depends only on the underlying network structure, i.e., on the preferential attachment parameter $m$. Since $m$ determines the minimum node degree, a larger $m$ generally reduces neighbor correlations. Consequently, as $m$ increases, the optimal $\alpha$ also increases. For very large $m$, neighbors are almost independent, implying $\alpha\approx 1$. This behavior is illustrated in Fig. 5, which shows the optimal $\alpha$ computed from (24) for different values of $m$. Moreover, note that the specific value of $\alpha$ also differs depending on the generation method of the networks, see appendix B for a comparison between BA model and configuration model.
It should be noted that the above results—especially regarding the validity of the mMFE and the robustness of $\alpha$—are accurate only for sufficiently large networks. In the examples discussed here, we found that $N\geq 10^{4}$ is a reasonable threshold. For smaller $N$, the mMFE, like other mean-field approximations, loses accuracy, see Fig. S4 in the appendix.
@@ IV Comparison to classical approximations
Figure 6: The shares $c_{1}$ and weighted shares $\bar{c}_{1}$ estimated from simulations, in comparison to the mMFE, MFE, and PA. Model parameters from regime A. The left plot shows the trajectories, while the right plot shows the $L_{1}$-error (21) between the mMFE and $\bar{\bm{c}}$, as well as between MFE and $\bm{c}$, and PA and $\bm{c}$.
In this section we compare the mMFE, which is a (M1)(M-1)-dimensional approximation, to other standard low-dimensional moment-closure approximations: the MFE, which also has dimension (M1)(M-1), and the pair-approximation (PA), which has a dimension of (M1)+M(M1)(M-1)+M(M-1). The PA results from closing the moment equations at second order and therefore includes the number of edges between different opinions as additional variables compared to the MFE. A detailed derivation of the PA for the CNVM is omitted here for brevity but the general idea as well as the resulting equations are presented in appendix C. As Fig. 6 illustrates, the classical approximations (MFE and PA) fail to capture the system behavior on BA networks, whereas the mMFE incurs only a very small error. This is not surprising, since the classical approximations assume a homogeneous network structure. The well-known heterogeneous variants of the MFE and PA, or the approximate master equation, are better suited for scale-free networks [gleeson2011high, Porter2016, Gleeson2013, Pugliese2009]. However, as these resolve the opinion shares per degree, they involve substantially more variables than the mMFE. For instance, on a BA network with $N=10^{4}$ nodes and preferential attachment parameter $m=4$, there are on average about 100 different node degrees, implying that the heterogeneous MFE has 100 times more variables than our mMFE. The quenched MFE [Mata2013] even resolves the expected state per node such that the required number of variables is $N$ times larger than that of the mMFE. Hence, to compare reduced order models with the same degree of efficiency, we only report comparison between the mMFE and the homogeneous MFE and PA.
Needless to say, the above mentioned methods and similar approaches, such as the heterogeneous MFE or approximate master equation, resolve (and hence can potentially predict) more detailed information about the dynamics. In contrast, our results indicate that the degree-weighted shares together with the mMFE provide a parsimonious macroscopic description that captures the relevant collective behavior of the CNVM on certain scale-free networks. The robustness and resulting universality of the parameter $\alpha$ make the mMFE a lightweight and computationally efficient surrogate when only a macroscopic quantity is of interest.
@@ V Conclusion
In this paper, we derived the mMFE (5), a modified version of the classical mean-field equation (MFE) that describes the dynamics in terms of weighted shares and includes an average “correlation factor” $\alpha$. We then verified that this equation provides an accurate and efficient approximation of the continuous-time noisy voter model on scale-free networks generated by the Barabási–Albert or the configuration model. It outperforms the standard MFE and pair-approximation (PA), while remaining low-dimensional. We found that the approximation error depends on the underlying parameter regime; in particular, in regimes C and D—corresponding to (partially) noise-free dynamics and/or very dissimilar imitation rates—a time-constant correlation factor $\alpha$ does not exist, causing larger errors. Obviously, arbitrary scale-free networks that may exhibit strong structural inhomogeneities—for instance, networks consisting of only sparsely connected communities—would require additional variables to retain accuracy, such as separate weighted shares for each community. Low-dimensional approximations for these types of networks could be investigated in future work.
The inspiration for the mMFE (5) arose from findings in [Luecke2024], where a computational approach was proposed to identify forecastable variables of dynamical processes on networks. For such processes, reduced modeling is typically guided by theoretical insights—our approach therefore constitutes an exception. Along these lines, it would be highly desirable to gain a theoretical understanding of the observed superior behavior of the mMFE. Furthermore, this work also highlights the importance of data-driven approaches, since many of the insights presented here were supported by computational analysis. These methods could be utilized in future work to derive low-dimensional approximations for other systems that are not yet well understood analytically.
This work in particular could also be extended by refining the mMFE in parameter regimes where its accuracy is limited, for instance by allowing a time-dependent $\alpha$. Moreover, a modified version of the PA that operates on the degree-weighted shares $\bar{\bm{c}}$ could be explored.
@@ Appendix A Derivation of mMFE for more than two opinions
Recall that in the case of $M>2$ opinions, the degree-weighted shares $\bar{\bm{c}}\in[0,1]^{M}$ are defined via $\bar{\bm{c}}=(\bar{c}_{a})_{a\in[M]}$,
@@
$\frac{\text{d}}{\text{d}t}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}]$ $=\sum_{b\neq a}\mathbb{P}(x_{i}=b)\ \mathbb{E}\big[q_{b,a}^{i}\mid x_{i}=b\big]$ (29)
$\ -\sum_{b\neq a}\mathbb{P}(x_{i}=a)\ \mathbb{E}\big[q_{a,b}^{i}\mid x_{i}=a\big].$ (30)
@@
Appendix B Scale-free networks generated by the configuration model
Figure S1: Error analysis for the configuration model, cf. Fig. 1 in the main text.
In this section we verify that the mMFE (5) is also valid for scale-free networks generated by the configuration model [Bollobas1980], which can be used to generate a random graph with an arbitrary degree sequence $(d_{i})_{i\in[N]}$. The method works as follows: After assigning $d_{i}$ half-edges to each node $i$, two half-edges are repeatedly picked from all remaining half-edges, joined together to form an edge, and removed from the set of available half-edges, until all half-edges have been used up. This potentially yields a multigraph since self-loops and multiple edges between two nodes are possible. In such a case we simply restart and repeat the whole procedure until finding a simple graph. To produce a scale-free network, we first generate a degree sequence $(d_{i})_{i\in[N]}$ such that the associated degree distribution follows a power law with exponent $\gamma=3$, and then apply the configuration model. To facilitate a fair comparison to a BA model with preferential attachment parameter $m$, the minimum degree in the degree sequence should also be chosen as $m$.
Asymptotically, the BA model and the configuration model are relatively similar. For example, for both models the average clustering coefficient falls to zero rapidly at a rate of approximately $1/N$, and the diameter grows moderately at a rate of approximately $\log N$ [Luecke2024a]. Furthermore, the assortativity, i.e., the degree correlation between neighbors, is asymptotically zero for both models. However, for the finite network sizes of $N\approx 10^{4}$ considered here, the BA model and the configuration model exhibit significant differences. The most noteworthy aspect is that for the BA model the assortativity is still clearly negative, meaning that it is more likely for small degree nodes to be connected to large degree nodes, while for the configuration model the assortativity is already close to zero. Moreover, the BA model shows a considerably larger clustering coefficient than the configuration model.
Results. We conducted the same error and robustness analysis for the configuration model as presented for the BA model in the main text. The error analysis in the four regimes A, B, C, D is shown in Fig. S1 and is almost identical to the BA model. The only significant difference is that for the configuration model the errors in the noise-less regime $C$ are lower. This is likely caused by the aforementioned lack of clustering and assortativity compared to the BA model, which hinders the formation of clusters of identical opinion.
The robustness analysis for the configuration model is shown in Fig. S2 and is also very similar to the BA model. The maximum error and the variance of errors is slightly larger for the configuration model, but the median error is also below 0.003.
The optimal correlation factor for the configuration model (with minimum degree 3) is $\alpha=0.75$, which is slightly smaller than the factor $\alpha=0.81$ determined for the BA model with preferential attachment parameter $m=3$ in the main text. This difference could be explained by the negative assortativity of the BA model which implies that small degree nodes are more likely connected to large degree nodes. Since such a pair of nodes with small and large degree is less correlated than a pair of nodes with small degrees, the correlation factor $\alpha$ is smaller for the configuration model (indicating a larger positive correlation of neighbors). A more thorough investigation could be conducted in future work.
Figure S2: Robustness of the mMFE for the configuration model, cf. Fig. 4 in the main text.
Appendix C Pair-approximation
In this section we briefly outline the derivation of the classical pair-approximation in the case of $M=2$ opinions. A more detailed derivation can readily be found in the literature, see for instance [Peralta2018, Luecke2024a]. Instead of closing the first-order moment equation using an independence assumption $\mathbb{E}[(1-x_{i})x_{j}]\approx\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}]$, the pair approximation assumes that all neighbor pairs are statistically equivalent and introduces the average pair probability $\mathbb{E}[(1-x_{i})x_{j}]=:s$ as an additional variable. Based on (II), this yields
$\frac{\text{d}}{\text{d}t}c=s(r_{0,1}-r_{1,0})-c(\tilde{r}_{0,1}+\tilde{r}_{1,0})+\tilde{r}_{0,1}.$ (37)
To obtain the evolution equation of $s$, we then have to consider the second-order moment equation, which in turn contains third-order moments, e.g., terms of the form $\mathbb{E}[x_{k}(1-x_{i})x_{j}]$ that describe the probability of a connected triplet of nodes kijk-i-j being in states $(1,0,1)$. We close this second-order moment equation by assuming that the outer two nodes $k$ and $j$ of the triplet are conditionally independent given $x_{i}$, which yields
$\mathbb{E}[x_{k}(1-x_{i})x_{j}]$ $\approx\frac{\mathbb{E}[(1-x_{i})x_{k}]}{1-\mathbb{E}[x_{i}]}\mathbb{E}[(1-x_{i})x_{j}]$ (38)
$\approx\frac{s^{2}}{1-c}.$ (39)
Following this procedure for all occurring third-order moments yields this two-dimensional system of equations called pair-approximation:
$\frac{\text{d}}{\text{d}t}c$ $=s(r_{0,1}-r_{1,0})-c(\tilde{r}_{0,1}+\tilde{r}_{1,0})+\tilde{r}_{0,1}.$ (40)
$\frac{\text{d}}{\text{d}t}s$ $=-2\frac{d-1}{d}\frac{s^{2}}{1-c}r_{0,1}-2\frac{d-1}{d}\frac{s^{2}}{c}r_{1,0}$ (41)
$\quad+s\Big(\frac{d-2}{d}r_{0,1}+\frac{d-2}{d}r_{1,0}-2\tilde{r}_{0,1}-2\tilde{r}_{1,0}\Big)$ (42)
$\quad+c\Big(\tilde{r}_{1,0}-\tilde{r}_{0,1}\Big)+\tilde{r}_{0,1},$ (43)
where $d$ denotes the average degree in the graph.
Appendix D Supplementary figures
Figure S3: Error of the mMFE for the parameters $r_{1,0}=1$, $\tilde{r}_{1,0}=\tilde{r}_{0,1}=0.01$ and varying $r_{0,1}$. The correlation factor $\alpha$ was optimized for each system, resulting in values between $0.8$ and $0.84$. The plot also shows the regions of regimes A and B as defined in the main text.
Figure S4: Error of the mMFE for the parameters $r_{0,1}=1$, $r_{1,0}=0.98$, $\tilde{r}_{0,1}=\tilde{r}_{1,0}=0.01$ and $\alpha=0.81$ against the number of nodes NN of the BA network.